# Reasoning or Reciting? Exploring the Capabilities and Limitations of Language Models Through Counterfactual Tasks

### v1

@@ Abstract
The impressive performance of recent language models across a wide range of tasks suggests that they possess a degree of abstract reasoning skills. Are these skills general and transferable, or specialized to specific tasks seen during pretraining? To disentangle these effects, we propose an evaluation framework based on “counterfactual” task variants that deviate from the default assumptions underlying standard tasks. Across a suite of 11 tasks, we observe nontrivial performance on the counterfactual variants, but nevertheless find that performance substantially and consistently degrades compared to the default conditions. This suggests that while current LMs may possess abstract task-solving skills to a degree, they often also rely on narrow, non-transferable procedures for task-solving. These results motivate a more careful interpretation of language model performance that teases apart these aspects of behavior.
{strip}
@@ 1 Introduction
We propose to measure such task-level generalizability by taking tasks on which LMs perform well, and altering the conditions or rules under which these tasks are performed. The general reasoning procedure for these tasks remains the same under the new conditions, but the specific input-output mappings are changed. We call the new tasks counterfactual tasks, as they deviate from the default, generally assumed conditions for these tasks. Figure 1 shows examples: in the top left, default arithmetic is performed in base-10, while counterfactual arithmetic is performed in base 9. If models implement a general and transferable task-solving procedure, we expect comparable performance on counterfactual and default tasks; if they employ procedures tailored to default task conditions, we expect a drop in the counterfactual performance.
@@
We evaluate the performance of GPT-4 (OpenAI, 2023), GPT-3.5, Claude (Anthropic, 2023), and PaLM-2 (Anil et al., 2023) on tasks under both the default and counterfactual conditions. We observe above-random counterfactual performance for most tasks, indicating some degree of task generalizability. However, their performance on counterfactual task variants consistently and substantially degrades relative to the performance on the default settings. This suggests that these models’ ability on these tasks is supported at least in part by non-transferable, default-condition-specific behaviors rather than abstract, generalizable reasoning skills.
These results also reveal several surprising relations between model behavior on default and counterfactual tasks (§5), including correlations between default and counterfactual performance, varying effectiveness of zero-shot chain-of-thought prompting Kojima et al. (2023), and interactions between task- and instance-level frequency effects. Overall, we find that small variations on the default instantiations of tasks are challenging for models, and thus the success of existing LMs should not be fully attributed to fully general capacity for the target task.
@@ 2 Counterfactual Tasks
where the $\operatorname*{arg\,max}$ is computed with an approximate decoding procedure and $\operatorname{prompt}_{f}$ and $\operatorname{prompt}_{w}$ are prompt templates that describe tasks and world models respectively. For each task, we devise one or more $w^{\text{cf}}$ that deviate from the default world (i.e., the default task conditions). We evaluate both $h(f,w^{\text{default}},x)$ and $h(f,w^{\text{cf}},x)$ via task-specific metrics. If we control $f_{w}(x)$ to be similarly hard for both $w^{\text{default}}$ and $w^{\text{cf}}$, we can attribute the performance difference to an LM overfitting to the default instantiation of the task.
@@ 3.1 Arithmetic
Modern LMs have been shown to possess basic numerical reasoning abilities (Lewkowycz et al., 2022), with even GPT-3 reporting near-perfect accuracy for two-digit additions (Brown et al., 2020). On the other hand, Razeghi et al. (2022) find that LMs perform significantly better on operations involving numbers that occur more frequently in the pretraining data, and Li et al. (2023d) show that symbol replacement affects the mathematical ability of BERT (Devlin et al., 2019)-like models; both findings point to overfitting and memorization effects. We consider the same two-digit addition task, the simplest arithmetic task in Brown et al. (2020), but inspect a model’s accuracy in different bases. We use base-8, 9, 11, and 16 as the counterfactual setup which are natural generalizations to base-10 arithmetic. These bases were chosen to control for task difficulty (see §7.1 for a discussion) and also to test for how relatively uncommon (9 & 11) and common (8 & 16) bases affect performance (see §5.1 for an analysis). To ensure the model understands the different bases, the CCC evaluates the successor relation under each base.
@@ 3.5 Spatial Reasoning
A major debate around LMs is whether grounded representations of meaning can be learned from form alone (Bender and Koller, 2020; Piantadosi and Hill, 2022; Mollo and Millière, 2023). Studies have shown that LMs can learn meaningful world representations through text-only training (Abdou et al., 2021; Li et al., 2023c; Jin and Rinard, 2023). In particular, Patel and Pavlick (2022) find that LMs learn representations of cardinal directions that can be aligned to grounded conceptual spaces with few-shot demonstrations.
@@ 3.6 Drawing
Despite being trained on only textual data, LMs have been shown to be able to structure their representations of perceptual concepts such as size and color (Abdou et al., 2021; Patel and Pavlick, 2022; Zhang et al., 2020; Ilharco et al., 2021; i.a.) in a way that credibly mirrors the physical world. Recent LMs can even generate plausible drawings of objects using code such as TikZ and SVG Bubeck et al. (2023); Zhang et al. (2023c). We evaluate the visual understanding of LMs by asking them to generate code for drawing various objects in the Processing language. Psychological studies have shown that humans have the ability to rotate mental representations of objects Shepard and Metzler (1971); Vandenberg and Kuse (1978). For the counterfactual settings, we similarly ask the LM to generate code that draws the same object, but rotated or vertically flipped. We disallow the use of functions such as rotate to prevent shortcut solutions (see §7.2 for further discussion). As with the spatial reasoning task (§3.5), an ideal model should be robust to these settings. For the CCC, we ask the model to draw a straight line at the top of the canvas in addition to the object; a flipped/rotated line thus signifies an understanding of the transformations.
@@ 3.9 SET Game
In each round, a player finds a SET of 3 cards in a 12-card board whose values for each attribute are either all the same or all unique. This game has been thoroughly studied in computer science, from the perspective of coding theory and combinatorics (Davis and Maclagan, 2003), linear algebra (Coleman and Hartshorn, 2012), and complexity theory (Chaudhuri et al., 2003). We suspect this popularity makes it susceptible to overfitting by LMs and investigate this possibility. We ask the LM to identify the card on a board that completes a 3-card SET with two given cards. In the counterfactual setup, we invert the rule for the number attribute, requiring its value to be mixed, in other words, neither all the same nor all unique. For the CCC, we ask the model for the validty of a SET under the original rule and the counterfactual rule.
@@ 4 Results
For each task, we evaluate GPT-4 (gpt-4-0314; OpenAI, 2023), GPT-3.5 (gpt-3.5-turbo-0301), Claude (claude-v1.3; Anthropic, 2023), and PaLM-2 (text-bison-001; Anil et al., 2023). As these are closed-source models, we do not have any information regarding their size, architecture, and pretaining details. [^6] We also explored open-source models in preliminary experiments, but found that they possess unsatisfactory instruction-following ability, to the point that often their output cannot be meaningfully parsed into a prediction. We therefore do not include these models. We note that the largest PaLM model is not publicly accessible, and we can only test the second-largest version. For each task, we experiment with both with and without encouraging the model to reason step by step, by adding the phrase “Let’s think step by step.” in our prompts (Kojima et al., 2023; Reynolds and McDonell, 2021). Following Kojima et al. (2023), we refer to this step-by-step setup as zero-shot chain-of-thought prompting (0-CoT; Nye et al., 2021; Wei et al., 2022). We include all prompts in §B.
Figures 2 and 3 show our results. §C contains the numeric version. We see a consistent pattern where LMs perform substantially worse on the counterfactual task variants, both with and without 0-shot CoT. For most cases, LMs exhibit an above-random counterfactual performance, suggesting some degree of the targeted ability. However, when the CCC accuracy is high, usually the case for GPT-4 and in select settings for other models too, the default-counterfactual gaps demonstrate limitations in the abstract capacity to solve the target task. When the CCC accuracy is lower, the failure of counterfactual world comprehension would be a confounder to this conclusion, but often the gaps are so large (sometimes even dropping from near-perfect to near-zero, such as for arithmetic) that they are nonetheless strongly indicative of non-transferable, default condition-specific implementations of the original task. The fact that the LMs sometimes cannot evaluate the CCC well under the counterfactual conditions, but can do so under the default conditions (e.g., for arithmetic, programming, drawing, etc.) itself also points to overfitting to the latter.
@@ 5 Analysis
Our counterfactual worlds are not designed to be completely alien to the LMs but only less common than the assumed default case. In this sense, the counterfactual-ness of these worlds is relative, and here we take a more nuanced look at how the commonness of these counterfactual conditions affects the default-counterfactual performance gap. For example, in the arithmetic task, all models perform better in bases 8 and 16, likely due to their relative abundance compared to bases 9 and 11. In spatial reasoning, the smallest counterfactual performance degradation is usually from when the north and south directions are swapped—even exceeding the default task performance for PaLM-2—potentially because some programming libraries use an inverted $y$-axis, such as matplotlib (Python), ggplot (R), and D3 (JavaScript) (see §A.5). For chord fingering, the common alternative drop-D tuning of guitars leads to the highest counterfactual performance. These correlations between the counterfactual performance and the commonness of the counterfactual worlds paint a more fine-grained picture than a binary default versus counterfactual distinction and point to a memorization-like effect where the models perform better under more common conditions.
@@
Figure 4: Investigating the relationship between the default task performance and counterfactual performance, broken down by different factors. Only GPT-4 with 0-shot CoT results are shown. There is a consistent default-counterfactual correlation across tasks and when varying different factors.
@@ 5.2 Proximity between Default and Counterfactual Conditions
Another axis along which the counterfactual worlds differ is in their proximity to the default conditions. For example, for the different arithmetic bases, bases 9 and 11 are closer to base 10, but less common than bases 8 and 16. While the default-counterfactual gap is most affected by commonness for the arithmetic task, for the guitar and ukulele tunings (other than the drop-D tuning), the LM performance generally decreases monotonically with the distance from the original tunings.
The FOLIO dataset (Han et al., 2022) enables another analysis of how proximity to the default conditions affects the model performance, without performing counterfactual perturbations. This dataset was constructed to mostly follow common sense, i.e., containing premises and conclusions that are deemed true in the real world. However, this is not always the case, with premises such as “John can make meals which are popular at the party,” whose factuality cannot be determined alone.
We evaluate how the distance between the world state described by the premises and the belief state of LMs influences LM performance by training a predictive model given features approximating this distance. For each test instance, we ask the LMs whether the premises and conclusion are true, false, or uncertain. We train a logistic regression model to predict LM correctness on each test instance, using as features the total number of premises in an input, the proportion of the premises that are true/false/uncertain, as encoded by the LM, as well as whether the LM-predicted truthfulness of the conclusion matches the label of the instance (that is, a feature that predicts the entailment/neutral/contradiction label of the instance from the truthfulness of the conclusion alone, ignoring premises).
Figure 5 shows the learned coefficients of these features, as well as their 95% confidence interval bootstrapping with 1,000 iterations (Efron and Tibshirani, 1993). Ideally, a robust model should predict solely based on symbolic deduction and extralinguistic truthfulness information should not affect its accuracy. In other words, these features should all have coefficients 0 and have no predictive power with respect to the model’s correctness. However, all LMs predict more correctly with more realistic (true) premises, and when the conclusion’s LM-predicted truthfulness matches the label. On the other hand, they tend to perform worse when there are more false or uncertain premises. Most of these trends are statistically significant. This means that the reasoning ability of LMs is affected by the distance between the real world (as believed by the LMs) which is the default condition and the world state under which reasoning is required.
@@
Recalling our formalization $h_{\text{LM}}(f,w,x)$ in §2, the previous two subsections analyzed how the commonness of $w$ and its proximity to $w^{\text{default}}$ affect the observed patterns. We now explore how the counterfactual performance correlates with the default task performance by varying the other three elements in this formalization: the task $f$, the input $x$, and the LM.
We first consider different task variants. For arithmetic, beyond 2-digit addition, we also measure GPT-4’s 3- and 4-digit addition performance (Figure 3(a)). [^7] See Dziri et al. (2023) for a related analysis but for multiplications. For note retrieval from melodies, we use the index of the inquired note as the proxy for difficulty (Figure 3(b)). [^8] This is not a new task variant as compared to the setup in §3.7, but rather a decomposition of our original results. For SET, while our original task shows two cards and asks a model to find the missing one from a 3-card SET, we change the task to instead only show one or none of the cards in a SET, while still requiring the model to identify the SET on a board (Figure 3(c)). For all these task variants, we see a strong correlation between the original and the counterfactual world performance.
We also see this effect when breaking down results by test instances. In Figure 3(d), we separate the different chord types, and observe that the default task performance correlates with the counterfactual performance. Similarly, reexamining our main results in Figures 2 and 3, for most tasks, stronger models under default conditions are also stronger models under counterfactual conditions, and vice versa. Overall, these correlations mean that the default task performance can be a good indicator of its counterfactual performance, and hence we should not discount the utility of traditional evaluations. [^9] Though these correlations are not necessarily causal.

### v3

@@ Abstract
The impressive performance of recent language models across a wide range of tasks suggests that they possess a degree of abstract reasoning skills. Are these skills general and transferable, or specialized to specific tasks seen during pretraining? To disentangle these effects, we propose an evaluation framework based on “counterfactual” task variants that deviate from the default assumptions underlying standard tasks. Across a suite of 11 tasks, we observe nontrivial performance on the counterfactual variants, but nevertheless find that performance substantially and consistently degrades compared to the default conditions. This suggests that while current LMs may possess abstract task-solving skills to an extent, they often also rely on narrow, non-transferable procedures for task-solving. These results motivate a more careful interpretation of language model performance that teases apart these aspects of behavior.
{strip}[Uncaptioned image]
@@ 1 Introduction
We propose to measure such task-level generalizability by taking tasks on which LMs perform well, and altering the conditions or rules under which these tasks are performed. The general reasoning procedure for these tasks remains the same under the new conditions, but the specific input-output mapping functions are changed. We call the new tasks counterfactual tasks, as they deviate from the default, generally assumed conditions for these tasks. Figure 1 shows examples: in the top left, default arithmetic is performed in base-10, while counterfactual arithmetic is performed in base 9. If models implement a general and transferable task-solving procedure, we expect comparable performance on counterfactual and default tasks; if they employ procedures tailored to default task conditions, we expect a drop in the counterfactual performance.
@@
We evaluate the performance of GPT-4 (OpenAI, 2023), GPT-3.5, Claude (Anthropic, 2023), and PaLM-2 (Anil et al., 2023) on tasks under both the default and counterfactual conditions. We observe above-random counterfactual performance for most tasks, indicating some degree of task generalizability. However, the performance on counterfactual task variants consistently and substantially degrades relative to the performance on the default settings. This suggests that these models’ ability on these tasks is supported at least in part by non-transferable, default-condition-specific behaviors rather than abstract, generalizable reasoning skills.
These results also reveal several surprising relations between model behavior on default and counterfactual tasks (§5), including correlations between default and counterfactual performance, varying effectiveness of zero-shot chain-of-thought prompting Kojima et al. (2023), and interactions between task- and instance-level frequency effects. Overall, we find that small variations on the default instantiations of tasks are challenging for models, and thus the success of existing LMs on standard benchmarks should not be considered as sufficient evidence for their possession of full general capacity for the target task.
@@ 2 Counterfactual Tasks
where the $\operatorname*{arg\,max}$ is computed with an approximate decoding procedure and $\operatorname{prompt}_{f}$ and $\operatorname{prompt}_{w}$ are prompt templates that describe tasks and world models respectively. For each task, we devise one or more $w^{\text{cf}}$ that deviate from the default world (i.e., the default task conditions). We evaluate both $h(f,w^{\text{default}},x)$ and $h(f,w^{\text{cf}},x)$ via task-specific metrics. If we control $f_{w}(x)$ to be similarly hard between $w^{\text{default}}$ and $w^{\text{cf}}$, we can attribute the performance difference to an LM overfitting to the default instantiation of the task.
@@ 3.1 Arithmetic
Modern LMs have been shown to possess basic numerical reasoning abilities (Lewkowycz et al., 2022), with Brown et al. (2020) even reporting near-perfect GPT-3 accuracy for two-digit additions. On the other hand, Razeghi et al. (2022) find that LMs perform significantly better on operations involving numbers that occur more frequently in the pretraining data, and Li et al. (2023d) show that symbol replacement affects the mathematical ability of BERT (Devlin et al., 2019)-like models; both findings point to overfitting and memorization effects. We consider the same two-digit addition task, the simplest arithmetic task in Brown et al. (2020), but inspect a model’s accuracy in different bases. We use base-8, 9, 11, and 16 as the counterfactual setup which are natural generalizations to base-10 arithmetic. These bases were chosen to control for task difficulty (see §7.1 for a discussion) and also to test for how relatively uncommon (9 & 11) and common (8 & 16) bases affect performance (see §5.1 for an analysis). To ensure the model understands the different bases, the CCC evaluates the successor relation under each base.
@@ 3.5 Spatial Reasoning
A major debate around LMs is whether grounded representations of meaning can be learned from form alone (Bender and Koller, 2020; Piantadosi and Hill, 2022; Mollo and Millière, 2023). Studies have shown that LMs can learn meaningful world representations through text-only training (Abdou et al., 2021; Li et al., 2023c; Jin and Rinard, 2023). In particular, Patel and Pavlick (2022) find that LMs learn representations of spatial relations and cardinal directions that can be aligned to grounded conceptual spaces with few-shot demonstrations.
@@ 3.6 Drawing
Despite being trained on only textual data, LMs have been shown to be able to structure their representations of perceptual concepts such as size and color (Abdou et al., 2021; Patel and Pavlick, 2022; Zhang et al., 2020; Ilharco et al., 2021; i.a.) in a way that credibly mirrors the physical world. Recent LMs can even generate plausible drawings of objects using code such as TikZ and SVG Bubeck et al. (2023); Zhang et al. (2023c). We evaluate the visual understanding of LMs by asking them to generate code for drawing various objects in the Processing language, which Sharma et al. (2024) found the LMs to be more adept in. Psychological studies have shown that humans have the ability to rotate mental representations of objects Shepard and Metzler (1971); Vandenberg and Kuse (1978). For the counterfactual settings, we similarly ask the LM to generate code that draws the same object, but rotated or vertically flipped. We disallow the use of functions such as rotate to prevent shortcut solutions (see §7.2 for further discussion). As with the spatial reasoning task (§3.5), an ideal model should be robust to these settings. For the CCC, we ask the model to draw a straight line at the top of the canvas in addition to the object; a flipped/rotated line thus signifies an understanding of the transformations.
@@ 3.9 SET Game
In each round, a player finds a SET of 3 cards in a 12-card board whose values for each attribute are either all the same or all unique. This game has been thoroughly studied in computer science, from the perspective of coding theory and combinatorics (Davis and Maclagan, 2003), linear algebra (Coleman and Hartshorn, 2012), and complexity theory (Chaudhuri et al., 2003). We suspect this popularity makes it susceptible to overfitting by LMs and investigate this possibility. We ask the LM to identify the card on a board that completes a 3-card SET with two given cards. In the counterfactual setup, we invert the rule for the number attribute, requiring its value to be mixed, in other words, neither all the same nor all unique. For the CCC, we ask the model for the validity of a SET under the original rule and the counterfactual rule.
@@ 4 Results
For each task, we evaluate GPT-4 (gpt-4-0314; OpenAI, 2023), GPT-3.5 (gpt-3.5-turbo-0301), Claude (claude-v1.3; Anthropic, 2023), and PaLM-2 (text-bison-001; Anil et al., 2023). As these are closed-source models, we do not have any information regarding their size, architecture, and pretaining details. [^6] We also explored open-source models in preliminary experiments, but found that they possess unsatisfactory instruction-following ability, to the point that often their output cannot be meaningfully parsed into a prediction. We therefore do not include these models. We note that the largest PaLM-2 model is not publicly accessible, and we can only test the second-largest version. For each task, we experiment both with and without encouraging the model to reason step by step, by adding the phrase “Let’s think step by step.” in our prompts (Kojima et al., 2023; Reynolds and McDonell, 2021). Following Kojima et al. (2023), we refer to this step-by-step setup as zero-shot chain-of-thought prompting (0-CoT; Nye et al., 2021; Wei et al., 2022). We include all prompts in §B.
Figures 2 and 3 show our results. §C contains the numeric version. We see a consistent pattern where LMs perform substantially worse on the counterfactual task variants, both with and without 0-shot CoT. For most cases, LMs exhibit an above-random counterfactual performance, suggesting some degree of the targeted ability. However, when the CCC accuracy is high, as is usually the case for GPT-4 and in select settings for other models too, the gaps in default vs. counterfactual task performance demonstrate limitations in their abstract capacity to solve the target task. When the CCC accuracy is lower, the failure of counterfactual world comprehension would be a confounder to this conclusion, but often the gaps are so large (sometimes even dropping from near-perfect to near-zero, such as for arithmetic) that they are nonetheless strongly indicative of non-transferable, default-condition-specific implementations of the original task. The fact that the LMs sometimes cannot evaluate the CCC well under the counterfactual conditions, but can do so under the default conditions (e.g., for arithmetic, programming, drawing, etc.) itself also points to overfitting to the latter.
@@ 5 Analysis
Our counterfactual worlds are not designed to be completely alien to the LMs but only less common than the assumed default case. In this sense, the counterfactual-ness of these worlds is relative, and here we take a more nuanced look at how the commonness of these counterfactual conditions affects the default-counterfactual performance gap. For example, in the arithmetic task, all models perform better in bases 8 and 16, likely due to their relative abundance compared to bases 9 and 11. In spatial reasoning, the smallest counterfactual performance degradation is usually from when the north and south directions are swapped—even exceeding the default task performance for PaLM-2—potentially because some programming libraries use an inverted $y$-axis, such as matplotlib (Python), ggplot (R), and D3 (JavaScript) (see §A.5). For chord fingering, the common alternative drop-D tuning of guitars (DADGBE) leads to the highest counterfactual performance for GPT-4. These correlations between the counterfactual performance and the commonness of the counterfactual worlds paint a more fine-grained picture than a binary default versus counterfactual distinction and point to a memorization-like effect where the models perform better under more common conditions.
@@
Figure 4: Investigating the relationship between the default task performance and counterfactual performance, broken down by different factors. Only GPT-4 with 0-shot CoT results are shown. There is a consistent default-counterfactual correlation across task variants when varying different factors.
@@ 5.2 Proximity between Default and Counterfactual Conditions
Another axis along which the counterfactual worlds differ is in their proximity to the default conditions. For example, for the different arithmetic bases, bases 9 and 11 are closer to base 10, but less common than bases 8 and 16. While the default-counterfactual gap is most affected by commonness for the arithmetic task, for the guitar and ukulele tunings (other than the drop-D tuning), the LM performance generally decreases monotonically with increasing distance from the original tunings.
The FOLIO dataset (Han et al., 2022) enables another analysis of how proximity to the default conditions affects LM performance, without counterfactual perturbations. This dataset was constructed to mostly follow common sense, with premises and conclusions deemed true in the real world. But this is not always the case, with premises like “John can make meals which are popular at the party,” whose factuality cannot be determined alone.
We evaluate how the distance between the (LM-believed) real world and the world state described by the premises (occasionally counterfactual to the LM) influences the LM’s performance by training a predictive model given features approximating this distance. For each test instance, we ask the LMs whether the premises and conclusion are true, false, or uncertain. We train a logistic regression model to predict LM correctness on each test instance, using as features the total number of premises in an input, the proportion of the premises that are true/false/uncertain, as encoded by the LM, as well as whether the LM-predicted truthfulness of the conclusion matches the label of the instance.
Figure 5 shows the learned coefficients of these features, as well as their 95% confidence interval bootstrapping with 1,000 iterations (Efron and Tibshirani, 1993). Ideally, a robust model should predict solely based on symbolic deduction and extralinguistic truthfulness information should not affect its accuracy. In other words, these features should all have coefficients 0 and have no predictive power with respect to the model’s correctness. However, all LMs predict more correctly with more realistic (true) premises, and when the conclusion’s LM-predicted truthfulness matches the label (indicating a tendency to predict the label solely based on the conclusion, ignoring premises). On the other hand, they perform worse when there are more false or uncertain premises. Most of these trends are statistically significant. This means that the reasoning ability of LMs is affected by the distance between the (LM-believed) real world and the world state under which the LMs are expected to reason.
@@
Recalling our formalization $h_{\text{LM}}(f,w,x)$ in §2, the previous two subsections analyzed how the commonness of $w$ and its proximity to $w^{\text{default}}$ affect the observed patterns. We now explore how the counterfactual performance correlates with the default task performance by varying the other three elements: the task $f$, the input $x$, and the LM.
We first consider different task variants with various difficulties. For arithmetic, beyond 2-digit addition, we also measure GPT-4’s 3- and 4-digit addition performance (Figure 4(a)). [^7] See Dziri et al. (2023) for a related analysis but for multiplications. For note retrieval from melodies, we use the index of the inquired note as the proxy for difficulty (Figure 4(b)). [^8] This is not a new task variant as compared to the setup in §3.7, but rather a decomposition of our original results. For SET, while our original task shows two cards and asks a model to find the missing one from a 3-card SET, we change the task to instead show one or none of the cards in a SET, while still requiring the model to identify the SET (Figure 4(c)). For all these task variants, we see a strong correlation between the original and counterfactual world performance.
We also see this effect when breaking down results by test instances $x$. In Figure 4(d), we separate the chord types, and observe that the default task performance correlates with the counterfactual performance. Similarly, reexamining our main results in Figures 2 and 3, for most tasks, stronger models under default conditions are also stronger models under counterfactual conditions, and vice versa. Overall, these correlations mean that the default task performance can be a good indicator of its counterfactual performance, and hence we should not discount the utility of traditional evaluations. [^9] Though these correlations are not necessarily causal.